OpenStudy (anonymous):

factor and solve 3p +1 = p^2 +3p^3

OpenStudy (xapproachesinfinity):

start with factoring p^2 in the left side \[3p+1=p^2+3p^3 \Longrightarrow 3p+1=(3p+1)p\]

OpenStudy (xapproachesinfinity):

now try it yourself

OpenStudy (xapproachesinfinity):

should be p^2 out side not p typo

OpenStudy (xapproachesinfinity):

\[3p+1=p^2+3p^3 \Longrightarrow 3p+1=(3p+1)p^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how am i supposed to get an answer from that